OpenStudy (anonymous):

What is the slope of a line parallel to the graph of the equation -x+2y = 4 ?

jimthompson5910 (jim_thompson5910):

first you must solve -x+2y = 4 for y what do you get when you do this?

OpenStudy (anonymous):

y= 1/2x+2

jimthompson5910 (jim_thompson5910):

good, this is in y = mx+b form

jimthompson5910 (jim_thompson5910):

where m = 1/2 b = 2

jimthompson5910 (jim_thompson5910):

m is the slope, so because m = 1/2, the slope is 1/2

jimthompson5910 (jim_thompson5910):

The slope of ANY line parallel to this one is equal to 1/2 (since parallel lines have equal slopes)

OpenStudy (anonymous):

so the answer is 1/2?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

ok, that was easy. haha. Thanks for the help

jimthompson5910 (jim_thompson5910):

sure thing, glad it makes sense